# Dark Energy Survey Year 3 results: Simulation-based wwCDM inference from weak lensing and galaxy clustering maps with deep learning. I. Analysis design

### v1

@@ III.1.2 Configuration
$N_{\textrm{cosmos}}$ $N_{\textrm{perts}}$ $L_{\textrm{box}}$ $N_{\textrm{part}}$
$N_{z\textrm{-shells}}$
@@
base 1 20 900 $832^{3}$ 140
$L_{\mathrm{box}}$ 1 20 2 2502\,250 $2\,080^{3}$ 140
$N_{\textrm{part}}$ 1 20 900 $2\,048^{3}$ 140
# $z$-shells 1 20 900 $832^{3}$ 500
@@ III.1.3 Baryonic Feedback
The dark matter particles in PkdGrav3 evolve under gravity alone. However, it has been shown that baryonic feedback can bias cosmological constraints, particularly on small scales [e.g. 84, 85, 86, 87].
@@ III.2 Map Projection
Following [38, 40, 8], we define the probe kernels of the weak lensing (WL) and intrinsic alignment (IA, see Sec. III.4.3) signal as
@@ III.2.3 Validation
The power spectra of the noiseless full-sky probe maps from the CosmoGridV1 have been validated against theoretical predictions from two independent codes. In [F22], comparison was performed with the PyCCL package 111111 CCL described in [105] for eqs. 2 and 3 using KiDS-1000 redshift distributions. In [K23], validation was conducted against the PyCosmo package  [^12] https://cosmology.ethz.ch/research/software-lab/PyCosmo.html introduced in [107] for eqs. 2, 3 and 4 using generic Stage-III-like redshift distributions. Both studies showed agreement within $5\%$, consistent with the minimal error expected from discrepancies between different theory predictions [108, 109].
@@ IV.1.1 Optimization Objective
between the output of the trainable compression $s\coloneqq S_{\varphi}(x)$ and the parameters $\theta$ is maximized, where $D_{KL}$ is the Kullback-Leibler (KL) divergence [128], and $H$ the Shannon entropy [129]. This expression is intractable because the exact posterior $p(\theta|s)$ is unknown. However, there exist a number of tractable lower bounds to eq. 14 that can instead serve as optimization targets [see e.g. 130]. Following [41, 48], and based on our empirical finding that this approach outperforms alternative estimators, we employ the variational lower bound [131]
@@ IV.1.2 Map Level
Following [F22, 39], our fiducial choice of architecture is the TensorFlow  [^13] https://www.tensorflow.org/ implementation of the DeepSphere 141414 deepsphere graph convolutional neural network developed in [110, 111]. In this design, the pixels making up the intricate footprint in Fig. 4 are represented as the nodes of a sparse graph that includes edges between a fixed number of neighbors.
@@ V.2.1 Synthetic Source Galaxy Catalog
Based on this processed sample, we can now generate the mock weak lensing shear maps $\gamma_{\rm WL}$. For the shape noise component $\gamma_{\rm SN}$, we sample the intrinsic ellipticities $\gamma_{j,\rm SN}$ and weights $w_{j}$ from the DES Y3 shape noise catalog $\tilde{\gamma}_{\textsc{Metacal}}$, consistent with the procedure adopted for the CosmoGridV1 in Sec. III.4.2. We choose not to include intrinsic alignment effects into the mocks.
@@ VI.1 Scale Cuts
Particular approximations in the forward model detailed in Sec. III are expected to fail at sufficiently small scales, most notably the assumption of linear galaxy biasing in our galaxy clustering maps. Consequently, implementing appropriate scale cuts that discard parts of the data is essential to ensure unbiased parameter inference. This strategy is consistent with previous DES Y3 analyses, where small-scale information was excluded [e.g. 5, 68, 20, 159] to mitigate potential systematic biases from, for example, inadequate modeling of baryonic effects.
@@ VI.1.1 Recovery of Buzzard Cosmology
Figure 7: Cosmological parameter recovery for individual probes and their combination using an ensemble of $N=15$ Buzzard mocks $x_{i}$ with map-level compression $s_{i}\coloneqq s(x_{i})$. In all cases, the MAP estimate of the posterior product $\prod_{i=1}^{N}p(\theta|s_{i})$ defined in eq. 18 (solid blue contours) lies within $0.5\sigma$ of a single observation $p(\theta|s_{i})$ shifted to align with the ground truth (solid red contour). This result demonstrates that potential systematic biases from modeling differences between the CosmoGridV1 and Buzzard simulations remain below the statistical uncertainty. Faint red contours show individual realizations $x_{i}$, contour levels represent 68th and 95th percentiles, and the gray shaded region indicates where the prior $p(\theta)$ is zero.
As a first validation, we test whether the pipeline correctly recovers the known cosmology of the Buzzard simulations. Following [156], we evaluate the product of posteriors from the ensemble of mocks $x_{i}$:
$p(\theta|\{s(x_{i})\})\coloneqq\prod_{i=1}^{N}p(\theta|s(x_{i})),$ (18)
where $N=15$. This approach assumes statistical independence between the $x_{i}$ and simulates the constraining power that would arise from a survey of $N$-fold larger area, thereby reducing the impact of cosmic variance.
We present the results for weak lensing and galaxy clustering individually, and their combination, in Fig. 7. For each configuration, the maximum a posteriori (MAP) estimates of $p(\theta|\{s(x_{i})\})$ lie within $0.5\sigma$ of the true cosmological parameters $\theta=\{\Omega_{m},S_{8},w\}$ in the two-dimensional marginals. This demonstrates that potential systematic biases from the distributional mismatch between the CosmoGridV1 and Buzzard simulations are subdominant to statistical uncertainties. We therefore conclude that for the considered scale cuts, the inference pipeline passes the test.
@@ VI.2 Posterior Coverage Tests
We validate this posterior through empirical coverage tests that measure the calibration of the distribution. For a test observation $x_{\mathrm{test}}\sim p(x|\theta_{\mathrm{test}})$ drawn from a known point in parameter space $\theta_{\mathrm{test}}$, a well-calibrated posterior $p(\theta|x_{\mathrm{test}})$ should produce nominal $(1-\alpha)\in[0,1]$ credible regions containing $\theta_{\mathrm{test}}$ with frequency $(1-\alpha)$. Coverage tests therefore involve repeating the inference procedure across many test observations $x_{\mathrm{test}}$ to measure this frequency. Although Bayesian credible regions are not theoretically required to satisfy frequentist coverage properties, demonstrating this correspondence empirically indicates that our inference procedure accurately quantifies parameter uncertainty.
@@
The tests of accuracy with random points (TARP) [162] method samples random reference points $\theta_{\mathrm{ref}}$ in parameter space and computes the fraction of posterior samples falling within a ball centered at $\theta_{\mathrm{ref}}$ and extending to the true value $\theta_{\mathrm{test}}$, thereby defining a credible region around the reference point. Repeating this procedure across a large number of reference points and observations $x_{\mathrm{test}}$ similarly yields a calibration curve that can be compared to the ideal diagonal. We implement TARP using publicly available code 181818 tarp. While TARP provides both necessary and sufficient conditions for well-calibrated posteriors, we include HPD as a complementary diagnostic since we have found it to be more sensitive in most cases.
@@ VII.1.1 Weak Lensing
We demonstrate in Appendix C that the marginal posterior contours of $A_{\mathrm{IA}}$ are not centered on the true value of zero due to projection effects.
@@ VII.1.2 Galaxy Clustering
Since the bispectrum contribution remains statistically significant at our fiducial smoothing scale of 32 Mpc/h for galaxy clustering [Fig. 1 in 167], and our map-level compression captures the beyond-Gaussian information content of the bispectrum in principle, we hypothesize that a similar degeneracy-breaking mechanism underlies the improvements observed in our analysis. A quantitative assessment of this plausible analogy is beyond the scope of this work.
@@ Appendix B Additional Figures
Appendix C Intrinsic Alignment Parametrization
@@
We show the posterior densities for the per-bin values $A_{\mathrm{IA}}^{i}$ computed from eq. 8 in Fig. C.1, where $i$ indexes the four redshift bins of the Metacal source galaxy sample. These two-dimensional marginals demonstrate that the highest posterior probabilities lie close to the true value of zero, consistent with the absence of intrinsic alignment modeling in our Buzzard mocks. We therefore conclude that the shift from zero toward positive $A_{\mathrm{IA}}$ values in Figs. 11, 12 and 14 results from a projection effect arising from the power-law parametrization of $A_{\mathrm{IA}}^{i}$ in terms of $A_{\mathrm{IA}}$ and $\eta_{A_{\mathrm{IA}}}$}.

### v2

# Dark Energy Survey Year 3 results: Simulation-based wwCDM inference from weak lensing and galaxy clustering maps with deep learning: Analysis design

@@ III Forward Modeling
The implementation of the forward model is available on GitHub 999 multiprobe-simulation-forward-model.
@@ III.1.2 Configuration
$N_{\textrm{cosmos}}$ $N_{\textrm{perts}}$ $L_{\textrm{box}}$ $N_{\textrm{part}}$ $N_{z\textrm{-shells}}$
@@
base 1 20 900 $832^{3}$ 69
$L_{\mathrm{box}}$ 1 20 2 2502\,250 $2\,080^{3}$ 69
$N_{\textrm{part}}$ 1 20 900 $2\,048^{3}$ 69
# $z$-shells 1 20 900 $832^{3}$ 392
@@ III.1.3 Baryonic Feedback
The dark matter particles in PkdGrav3 evolve under gravity alone. However, it has been shown that baryonic feedback can bias cosmological constraints, particularly on small scales [e.g. 85, 86, 87, 88].
@@ III.2 Map Projection
Following [38, 40, 8], we define the probe kernels of the weak lensing (WL, first introduced in [103]) and intrinsic alignment (IA, see Sec. III.4.3) signal as
@@ III.2.3 Validation
The power spectra of the noiseless full-sky probe maps from the CosmoGridV1 have been validated against theoretical predictions from two independent codes. In [F22], comparison was performed with the PyCCL package 121212 CCL described in [107] for eqs. 2 and 3 using KiDS-1000 redshift distributions. In [K23], validation was conducted against the PyCosmo package  [^13] https://cosmology.ethz.ch/research/software-lab/PyCosmo.html introduced in [109] for eqs. 2, 3 and 4 using generic Stage-III-like redshift distributions. Both studies showed agreement within $5\%$, consistent with the minimal error expected from discrepancies between different theory predictions [110, 111].
@@ IV.1.1 Optimization Objective
between the output of the trainable compression $s\coloneqq S_{\varphi}(x)$ and the parameters $\theta$ is maximized, where $D_{KL}$ is the Kullback-Leibler (KL) divergence [131], and $H$ the Shannon entropy [132]. This expression is intractable because the exact posterior $p(\theta|s)$ is unknown. However, there exist a number of tractable lower bounds to eq. 14 that can instead serve as optimization targets [see e.g. 133]. Following [41, 48], and based on our empirical finding that this approach outperforms alternative estimators, we employ the variational lower bound [134]
@@ IV.1.2 Map Level
Following [F22, 39], our fiducial choice of architecture is the TensorFlow  [^15] https://www.tensorflow.org/ implementation of the DeepSphere 161616 deepsphere graph convolutional neural network developed in [112, 113]. In this design, the pixels making up the intricate footprint in Fig. 4 are represented as the nodes of a sparse graph that includes edges between a fixed number of neighbors.
@@ IV.2 Neural Likelihood Estimation
Code available on GitHub 202020 multiprobe-simulation-inference.
@@ V.2.1 Synthetic Source Galaxy Catalog
Based on this processed sample, we can now generate the mock weak lensing shear maps $\gamma_{\rm WL}$. For the shape noise component $\gamma_{\rm SN}$, we sample the intrinsic ellipticities $\gamma_{j,\rm SN}$ and weights $w_{j}$ from the DES Y3 shape noise catalog $\tilde{\gamma}_{\textsc{Metacal}}$, consistent with the procedure adopted for the CosmoGridV1 in Sec. III.4.2. We choose not to include intrinsic alignment effects into the mocks, equivalent to an intrinsic alignment amplitude $A_{\mathrm{IA}}$ of zero.
@@ VI.1 Scale Cuts
Particular approximations in the forward model detailed in Sec. III are expected to fail at sufficiently small scales, most notably the assumption of linear galaxy biasing in our galaxy clustering maps. Consequently, implementing appropriate scale cuts that discard parts of the data is essential to ensure unbiased parameter inference. This strategy is consistent with previous DES Y3 analyses, where small-scale information was excluded [e.g. 5, 68, 20, 163] to mitigate potential systematic biases from, for example, inadequate modeling of baryonic effects.
@@ VI.1.1 Recovery of Buzzard Cosmology
Figure 7: Cosmological parameter recovery for individual probes and their combination using an ensemble of $N=15$ Buzzard mocks $x_{i}$ with map-level compression $s_{i}\coloneqq s(x_{i})$. In all cases, the MAP estimate of the joint posterior $p(\theta|\{s_{i}\}_{i=1}^{N})$ defined in eq. 18 (solid blue contours) lies within $0.5\sigma$ of a single observation $p(\theta|s_{i})$ shifted to align with the ground truth (solid red contour). This result demonstrates that potential systematic biases from modeling differences between the CosmoGridV1 and Buzzard simulations remain below the statistical uncertainty. Faint red contours show individual realizations $x_{i}$, contour levels represent 68th and 95th percentiles, and the gray shaded region indicates where the prior $p(\theta)$ is zero.
As a first validation, we test whether the pipeline correctly recovers the known cosmology of the Buzzard simulations. Following [160], we evaluate the joint posterior over the ensemble of mocks $x_{i}$:
$p(\theta|\{s(x_{i})\}_{i=1}^{N})\propto p(\theta)\prod_{i=1}^{N}p(s(x_{i})|\theta),$ (18)
where $N=15$. This expansion assumes conditional independence between the $x_{i}$ given $\theta$ and simulates the constraining power of a survey with $N$-fold larger area, thereby reducing the impact of cosmic variance. The independence assumption is justified because, analogously to our CosmoGridV1 mocks, we remove potential large-scale correlations between mock observations constructed from disjoint sky regions of the same underlying $N$-body simulation by applying an $\ell_{\textrm{min}}$-cut as detailed in Sec. III.6.2. Furthermore, the respective noise realizations are independent by construction.
We present the results for weak lensing and galaxy clustering individually, and their combination, in Fig. 7. For each configuration, the maximum a posteriori (MAP) estimates of $p(\theta|\{s(x_{i})\}_{i=1}^{N})$ lie within $0.5\sigma$ of the true cosmological parameters $\theta=\{\Omega_{m},S_{8},w\}$ in the two-dimensional marginals. This demonstrates that potential systematic biases from the distributional mismatch between the CosmoGridV1 and Buzzard simulations are subdominant to statistical uncertainties. We therefore conclude that for the considered scale cuts, the inference pipeline passes the test.
@@ VI.2 Posterior Coverage Tests
We validate this posterior through empirical coverage tests that measure the calibration of the distribution using held-out simulations. Specifically, for a test observation $x_{\mathrm{test}}\sim p(x|\theta_{\mathrm{test}})$ drawn from a known point in parameter space $\theta_{\mathrm{test}}$, a well-calibrated posterior $p(\theta|x_{\mathrm{test}})$ should produce nominal $(1-\alpha)\in[0,1]$ credible regions containing $\theta_{\mathrm{test}}$ with frequency $(1-\alpha)$. Coverage tests therefore involve repeating the inference procedure across many test observations $x_{\mathrm{test}}$ to measure this frequency. Although Bayesian credible regions are not theoretically required to satisfy frequentist coverage properties, demonstrating this correspondence empirically indicates that our inference procedure accurately quantifies parameter uncertainty.
@@
The tests of accuracy with random points (TARP) [166] method samples random reference points $\theta_{\mathrm{ref}}$ in parameter space and computes the fraction of posterior samples falling within a ball centered at $\theta_{\mathrm{ref}}$ and extending to the true value $\theta_{\mathrm{test}}$, thereby defining a credible region around the reference point. Repeating this procedure across a large number of reference points and observations $x_{\mathrm{test}}$ similarly yields a calibration curve that can be compared to the ideal diagonal. We implement TARP using publicly available code 212121 tarp. While TARP provides both necessary and sufficient conditions for well-calibrated posteriors, we include HPD as a complementary diagnostic since we have found it to be more sensitive in most cases.
@@ VII.1.1 Weak Lensing
We demonstrate in Sec. C.1 that the marginal posterior contours of $A_{\mathrm{IA}}$ are not centered on the true value of zero due to projection effects. In Sec. C.2, we present posterior results on CosmoGridV1 internal mock observations with $A_{\mathrm{IA}}\neq 0$ to show accurate parameter recovery in the presence of intrinsic alignments.
@@ VII.1.2 Galaxy Clustering
Since the bispectrum contribution remains statistically significant at our fiducial smoothing scale of 32 Mpc/h for galaxy clustering [Fig. 1 in 171], and our map-level compression captures the beyond-Gaussian information content of the bispectrum in principle, we hypothesize that a similar degeneracy-breaking mechanism underlies the improvements observed in our analysis. A quantitative assessment of this plausible analogy is beyond the scope of this work.
@@ 49 Institute for Astronomy, University of Edinburgh, Edinburgh EH9 3HJ, UK
51 INAF-Osservatorio Astronomico di Trieste, via G. B. Tiepolo 11, I-34143 Trieste, Italy
@@ Appendix B Additional Figures
Appendix C Intrinsic Alignment
This appendix presents additional results regarding the source galaxy intrinsic alignments modeled according to Sec. III.4.3.
C.1 Projection Effects
@@
We show the posterior densities for the per-bin values $A_{\mathrm{IA}}^{i}$ computed from eq. 8 in Fig. C.1, where $i$ indexes the four redshift bins of the Metacal source galaxy sample. These two-dimensional marginals demonstrate that the highest posterior probabilities lie close to the true value of zero, consistent with the absence of intrinsic alignment modeling in our Buzzard mocks (see Sec. V.2.1). We therefore conclude that the shift from zero toward positive $A_{\mathrm{IA}}$ values in Figs. 11, 12 and 14 results from a projection effect arising from the power-law parametrization of $A_{\mathrm{IA}}^{i}$ in terms of $A_{\mathrm{IA}}$ and $\eta_{A_{\mathrm{IA}}}$}.
C.2 CosmoGridV1 Internal Mocks
Since we do not include intrinsic alignments in our Buzzard mocks, we compare different intrinsic alignment amplitudes $A_{\mathrm{IA}}$ using mock observations derived from our weak lensing CosmoGridV1 forward model detailed in Sec. III.4.
Figure C.2: Marginal posterior probability densities for CosmoGridV1 internal mocks of varying intrinsic alignment amplitude $A_{\mathrm{IA}}$. For the intrinsic alignment parameters, the axis limits mark the prior boundaries from Tab. 1. In all cases, the inference pipeline accurately recovers the true parameters.
We present results for $A_{\mathrm{IA}}\in\{-1,0,1\}$, ${\eta_{A_{\mathrm{IA}}}}=0$, and $b_{\mathrm{TA}}=0$ in Fig. C.2 and find that our inference pipeline accurately recovers the ground truth values in all cases. The parameters $\eta_{A_{\mathrm{IA}}}$} and $b_{\mathrm{TA}}$ are only weakly constrained and partially hit the prior boundaries. Everything besides $A_{\mathrm{IA}}$ being fixed, the marginal distributions of the cosmological parameters are only mildly affected.